# Floralens: A Deep Learning Model for the Portuguese Native Flora

### v2

@@ 5.3 PlantCLEF and Wikipedia Test Sets
We considered two additional test sets: a random sample of 10,000 labeled images from the PlantCLEF’22-23 [56, 43] competition, and a sample of close to 1,500 images from Wikipedia. The PlantCLEF data we used was only a small sample of the entire “trusted” training set of PlantCLEF [58] that comprises approximately 2.9 million images covering 80,000 plant species. The repository is trusted as the image labels were obtained from academic sources or collaborative platforms like Pl@ntNet or iNaturalist. Our subset was built by first filtering out species not covered by the Floralens model, obtaining data for 1,593 (out of 1,678) species, and then randomly sampling 10,000 images.
@@
As for the Wikipedia test set, the images were identified through the Wikimedia REST API search functionality [59]. For each species in the Floralens domain, we used the species name as the keyword for a REST API search. Among other information, the search result typically yields a reference to an image stored at Wikipedia which we then considered for addition to the test set. After obtaining the images, we filtered out illustrations, herbarium specimens, and duplicates (associated with more than one species). Duplicates typically arise because the search may yield an image of a different species in the same genus, e.g., if the target species’ name does not have a Wikipedia page. Through this process, we obtained a dataset of 1,351 images for an equal number of species (one image per species). Compared to the PlantCLEF test set, the identifications associated with these images are less reliable as they result from an uncontrolled crowd-sourced effort with no specific directives for validation.
@@
Table 4 shows the results of the Floralens model for these test sets considered in terms of the Top-1, Top-5, and MRR metrics. We also recall the overall FTLS results (from Table 3) for easy comparison. We can observe that the PlantCLEF and Wikipedia test set results are marginally lower than those obtained for the FLTS, (by 0.02/0.03 in all metrics). The results show that the Floralens model performs well with datasets other than our base test suite.
Table 4: Top-1, Top-5, and MRR of the Floralens model for all test sets.
Dataset #I #S Top-1 Top-5 MRR
FLTS 29,360 1,678 0.67 0.86 0.75
PlantCLEF 10,000 1,593 0.65 0.84 0.73
Wikipedia 1,351 1,351 0.65 0.84 0.72
5.4 Genus Identification Results
@@
Table 5 shows the results of genus (as opposed to species) identification. The classification score for a genus is obtained by summing the scores for all the species in that genus output by the model. In [14] we created a model specific for genus identification but found that the results were essentially the same as obtained using this aggregation method. The greatest enhancement is observed for the Wikipedia test set, especially, for the Top-1 result ($\Delta=+0.14$). This probably originates from our previous observation that even when the image on the page for a species is wrongly labeled, Wikipedia does manage to provide an image of a plant of the same genus. The improvements observed for FLTS and PlantCLEF for all metrics are the same.
Table 5: Top-1, Top-5, and MRR of Floralens for genus prediction ($\Delta$: variation relative to species results).
Dataset Top-1 $\Delta$ Top-5 $\Delta$ MRR $\Delta$
FLTS 0.76 +0.09 0.91 +0.05 0.82 +0.07
PlantCLEF 0.74 +0.09 0.89 +0.05 0.80 +0.07
Wikipedia 0.79 +0.14 0.91 +0.07 0.83 +0.08
@@ 5.5 Comparative Pl@ntNet API Results
We now provide results comparing the Floralens model with models accessible via the Pl@ntNet API [57]. The Pl@ntNet API is a web service providing access to the same visual identification models used by state-of-the-art Pl@ntNet apps [6]. The API lets us obtain a set of ranked species for a given image for two models for worldwide flora: a so-called “legacy” model from 2022 (henceforth ${\rm PN}^{22}$) generated using CNN, and; a recent model announced in July 2023 [19], generated using Vision Transformers, henceforth ${\rm PN}^{23}$. Through the API, it is also possible to filter results from the ${\rm PN}^{23}$ model so that only species occurring in a specific biogeographic region are included. One such region is Southwestern Europe including Portugal, allowing the most head-to-head comparison between Floralens and Pl@ntNet possible. These results are identified by ${\rm PN}^{23\text{F}}$.
Table 6: Pl@ntNet API: comparative MRR values ($\Delta$: variation relative to the Floralens model).
Dataset ${\rm PN}^{22}$ $\Delta$ ${\rm PN}^{23}$ $\Delta$ ${\rm PN}^{23\text{F}}$ $\Delta$
FLTS 0.68 --0.07 0.80 +0.05 0.80 +0.05
PlantCLEF 0.72 --0.01 0.79 +0.06 0.79 +0.06
Wikipedia 0.73 +0.01 0.78 +0.06 0.79 +0.07
Table 7: Floralens vs Pl@ntNet API: MRR per data source in the FLTS ($\Delta$: variation relative to the Floralens model).
Source ${\rm PN}^{22}$ $\Delta$ ${\rm PN}^{23}$ $\Delta$ ${\rm PN}^{23\text{F}}$ $\Delta$
FloraOn 0.58 --0.17 0.79 +0.04 0.79 +0.04
iNaturalist 0.67 --0.10 0.81 +0.04 0.81 +0.04
Observation.org 0.59 --0.13 0.76 +0.04 0.76 +0.04
Pl@ntNet 0.77 +0.03 0.84 +0.10 0.84 +0.10
[…]
FLTS 0.68 --0.07 0.80 +0.05 0.80 +0.05
Table 6 shows the variation ($\Delta$) of the MRR values obtained for ${\rm PN}^{22}$, ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$ for all the test sets relative to the corresponding values obtained for Floralens. ${\rm PN}^{22}$ performs worse than Floralens for the FLTS, a variation of $-0.07$. The MRR values of ${\rm PN}^{22}$ are otherwise similar for PlantCLEF ($-0.01$) and Wikipedia ($+0.01$). The discrepancy observed for FLTS merits further analysis and is discussed below. Focusing now on ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$, the MRR values across all test sets range from 0.78 to just 0.80, and perform better than Floralens by a factor of 0.05 to 0.07. The Southwestern Europe species filter associated with ${\rm PN}^{23\text{F}}$ has little impact on the results.
[…]
Overall, the Floralens results are on par and in some cases better than ${\rm PN}^{22}$, and marginally worse than ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$.
6 Software Artifacts
@@ 6.1 Biolens Website
The Floralens model has been integrated into the Biolens project website [9]. As illustrated in the screenshots of Figure 8, the functionality is quite simple: (a) users submit photos of interest, and (b) obtain suggestions for identifications, with a textual indication of the model’s confidence. For instance, a “HIGH” confidence label as shown in (b) means that the model has a score higher than 70% for the species at stake; in other cases, we can have “MEDIUM” or “LOW” confidence labels corresponding to a score of 40-70% or lower than 40%, respectively. Up to 5 species are listed by the app, as long as the model outputs a minimum score of 15% for each identification. The Biolens website is hosted by a small virtual machine that requires just 2 CPU cores and 8 GB of RAM. With this configuration, invoking the model takes an average of 900 milliseconds per image, as calculated directly from approximately 28,000 requests to the server between April 2021 and September 2024. As for the model itself, it takes only 8 MB of disk storage. The server configuration is lightweight as we use the TF Lite variant of the Floralens model running on the server side (similarly to the other models for other taxa also hosted on the site). The user’s browser merely displays the server’s results, it does not host the model.
@@
Figure 8: Biolens – web application screenshots.
@@ 6.2 Biolens App
Figure 9: Biolens – mobile application screenshots.
@@ 6.3 Dataset and Results
The full Floralens dataset is publicly available on Zenodo [16]. The dataset contains the mapping between the image labels (ground truth), the image URLs from which they were retrieved, URLs for a site we maintain where all images are also stored, and GBIF identifiers when applicable (all images except those obtained from FloraOn). Ground truth and URLs are also available for the PlantCLEF and Wikipedia datasets used in the evaluation of Section 5, as well as for all datasets the Top-5 results and corresponding confidence levels for the Floralens model and the three Pl@ntNet model variations.
7 Conclusions
@@
In this paper, we describe the construction of a dataset for the Portuguese flora and the derivation of a deep-learning model for the automatic identification of the species therein. The universe of species was taken from the FloraOn dataset, provided by the Sociedade Portuguesa de Botânica and compiled exclusively by specialists. The dataset was constructed based on high-quality data from several research-grade datasets available via GBIF. Besides FloraOn these include: iNaturalist, Pl@ntNet, and Observation.org. We made the dataset available to the community on Zenodo [16]. The Floralens model was derived from this dataset using GAMLV which provides users with tools to derive models from datasets using off-the-shelf convolutional deep neural networks. The model is available online at the BioLens Project website and can also be used as part of the Biolens mobile application (e.g., offline, in the field).
The Floralens model has good predictive power, with an AUC metric value of 0.72 and, for a reference confidence level of 0.5, values for precision and recall of 0.85 and 0.53, respectively. It also features a relatively homogeneous predictive power across all data sources used in the dataset, with a maximum variation of 0.06, and values for Top-1 and Top-5 of 0.67 and 0.86, respectively. Compared with the state-of-the-art platform Pl@ntnet, Floralens performed on par with the “legacy 2022” model and only marginally worse when compared with the most recent model.
As for future work, we aim to improve the species coverage and the model’s accuracy. One way to do that is to include data from other datasets such as those of Encyclopedia of Life [60] and FloraIncognita [5]. With hindsight, also including images from Morocco and Algeria would probably have helped enhance the dataset with images from species in common with the Mediterranean flora.
We also want to address limitations that arise from the inconsistent use of taxonomic names and synonyms. Our list of Portuguese native species taken from FloraOn is as complete and up-to-date as possible. Recent taxonomic revisions changed the binomial names for some species. This is not necessarily reflected in the public datasets. For example, the species featured in the FloraOn listing as Atractylis gummifera is now known as Chamaeleon gummifer. Although not ubiquitous, it is widespread in the Mediterranean region [61]. Nevertheless, it is not included in the Floralens dataset as there weren’t enough ($\geq$ 50) photos available in GBIF. However, a recent query for Chamaeleon gummifer yields more than enough photos to include the species in the Floralens dataset in a future update.
More work is also required on the Biolens mobile app to improve its usability and optimize resource usage. Integrating with existing Citizen Science platforms is one possibility, allowing the user to upload Biolens records.
Finally, we plan to continue preliminary work on the use of image similarity models [14]. These may provide an alternative way to clinch an identification when GAMLV-based models yield low-confidence results. Hybrid classification models that combine both approaches are an interesting possibility.

### v3

@@ 4.2 PlantCLEF and Wikipedia Test Sets
We considered two additional test sets: a random sample of 10,000 labeled images from the PlantCLEF’22-23 [59, 46] competition, and a sample of close to 1,500 images from Wikipedia. The PlantCLEF data we used was only a tiny sample of the entire “trusted” training set of PlantCLEF [60] that comprises approximately 2.9 million images covering 80,000 plant species. The repository is trusted as the image labels were obtained from academic sources or collaborative platforms like Pl@ntNet or iNaturalist. Our subset was built by first filtering out species not covered by the Floralens model, obtaining data for 1,593 (out of 1,678) species, and then randomly sampling 10,000 images.
@@
As for the Wikipedia test set, the images were identified through the Wikimedia REST API search functionality [61]. For each species in the Floralens domain, we used the species name as the keyword for a REST API search. Among other information, the search result typically yields a reference to an image stored at Wikipedia, which we then considered for addition to the test set. After obtaining the images, we filtered out illustrations, herbarium specimens, and duplicates (associated with more than one species). Duplicates typically arise because the search may yield an image of a different species in the same genus, e.g., if the target species’ name does not have a Wikipedia page. Through this process, we obtained a dataset of 1,351 images for an equal number of species (one image per species). Compared to the PlantCLEF test set, the identifications associated with these images are less reliable as they result from an uncontrolled crowd-sourced effort with no specific directives for validation.
@@
Figure 9(a) shows the results of the Floralens model for these test sets considered in terms of the Top-1, Top-5, and MRR metrics for species prediction. We also recall the overall FTLS results (from Figure 7(a)) for easy comparison. We can observe that the PlantCLEF and Wikipedia test set results are marginally lower than those obtained for the FLTS (0.02 to 0.03 in all metrics). The results suggest that the Floralens model performs well with datasets other than our base test suite.
4.3 Genus and Family Identification Results
(a) Species.
(b) Genus.
(c) Family.
Fig 9: Species, genus, and family identification results for all test datasets (a,b and c, respectively).
@@ 4.3 Genus and Family Identification Results
Figures 9(b) and 9(c) show the results of genus and family identification for all datasets. The classification score for a genus/family is obtained by summing the scores for all the species in that genus/family output by the model. In [14] we created a model specific for genus identification but found that the results were essentially the same as obtained using this aggregation method. Genus results outperform species results (in Figure 9(a)) by 0.09 to 0.14 for Top-1, 0.05 to 0.07 for Top-5, and 0.07 to 0.11 for MRR. In turn, the family results outperform genus results by 0.07 to 0.08 for Top-1, 0.04 to 0.05 for Top-5, and 0.06 to 0.08 for MRR. As with species, the genus and family results are homogenous across all datasets. The more significant variation is observed for the genus results of the Wikipedia test set, e.g., the Top-1 result is 0.14 higher than that for the species. This might be explained by the fact that even when the image on the page for a species is wrongly labeled, Wikipedia does manage to provide one for a plant of the same genus (and by implication of the same family).
4.4 Classification using Multiple Images
It is often the case that an observation of a plant in the wild has multiple images associated with it to enable better identification. In this section, we take advantage of this fact and compute an aggregate score for each such observation based on the scores returned by Floralens for each image of a given specimen. We apply this operation to all observations with two or more associated images in the FLTS.
Fig 10: Results for classification using multiple images.
We identified 1185 such observations, corresponding to a total of 2480 images. More precisely, most observations have 2 images, 1090 out of 1185 (91%), and the remaining 95 observations have 3 or more images (82 have 3 images, 12 have 4 images, and a standalone observation has 6 images). The observations at stake are grouped by the GBIF identifier of images in the iNaturalist, Observation.org, and Pl@ntNet datasets. We cannot determine which images are related to the same observation for images obtained from FloraOn. Our analysis was not extended to the PlantCLEF dataset for the same reason. In the Wikipedia dataset, there is only one image per species.
For the 1185 observations of single specimens, we calculated an aggregate classification score by taking the average of the scores of the individual images. Figure 10 compares the standalone image classification (the baseline FLTS results) vs. the aggregate classification of observations in terms of Top-1, Top-5, and MRR. We observe clear score improvements: $+0.13$ for Top-1 (0.67 to 0.80), $+0.08$ for Top-5 (0.86 to 0.94), and $+0.12$ for MRR (0.74 to 0.86). The figure also depicts (rightmost) the results for the (comparatively few) 95 observations (9% of the cases) with 3 or more images showing even more pronounced improvements.
4.5 Classification using Geographical Location
(a) CDF plot for the shortest distance between GBIF coordinates and FloraOn grid records.
(b) Top-1, Top-5, and MRR variations for different values of $D$ in the geographical location filter.
(c) Images with altered ground truth ranks for different values of $D$ in the geographical location filter
Fig 11: Results for geographic data filter.
Some automated classification platforms use the geographical location associated with observations to improve the accuracy of the results. For instance, Pl@ntNet provides a parameter for the geographical zone (e.g., Southwestern Europe) and filters out species that do not occur in that region from the classification results, regardless of the output of the image classification model. We analyzed the impact of this strategy by making use of two items of information: (1) detailed FloraOn records of the geographical location of observations in Continental Portugal (excluding the Azores and Madeira archipelagos), and (2) the geographical location associated with FLTS images obtained from GBIF.
FloraOn geographic distribution data is based on the Military Grid Reference System (MGRS) grid zone with a resolution of 10$\times$10 km. For any given species, it keeps a set of MGRS grid elements for which observations have been reported. We collected this data for almost all species covered by the Floralens dataset – 1654 out of 1678 (98%). Next, we identified all images related to observations in Continental Portugal in the FLTS through GBIF. These amount to 3172 images covering 770 species. We analyzed the concordance between geographical locations and GBIF, by calculating for each image the shortest distance between the GBIF coordinates and the FloraOn grid for the ground truth species at stake. The cumulative distribution function (CDF) of the shortest distances is shown in Figure 11(a). Note that the distances in the x-axis have a logarithmic scale. The results indicate good agreement between FloraOn and image coordinates: the 95-th, 97-th, and 99-th percentiles of the distribution of distances are 18.9 km, 24.7 km, and 42.4 km.
When we supplied the model with an image taken at a given location, it replied with a ranking of suggested species. For each species in the ranking, we calculated the distance $d$ of the observation’s location to the closest known observation of that species. We discarded all species from the ranking output by the model such that $d>D$, where $D$ is a parameter. We applied this geographical filter for varying values of the distance threshold $D$ from 20 to 200 km with increments of 10 km. Note that the granularity of the MGRS grid used in FloraOn is 10km, thus justifying the choice of 20 km as the lowest value for $D$. To gauge the impact of the filter, we calculated the variations in the Top-1, Top-5, and MRR metrics versus the respective values without applying the filter. With no filter, we obtain 0.74 for Top-1, 0.89 for Top-5, and 0.80 for MRR. The variations are shown in Figure 11(b). As illustrated, we observe small positive variations, allowing us to conclude that using a geographical filter is beneficial if not significantly so. The highest variations are in the case of $D=40$ km: $+0.04$ for Top-1, $+0.02$ for Top-5, and $+0.03$ for MRR. As $D$ grows larger, the impact of the filter becomes less significant. The operation of the filter is also illustrated in Figure 11(c) in terms of the fraction of images whose classification results have been affected by the filter. The plot distinguishes the following two cases: (1) the ground truth (species) rank is improved, and; (2) the ground truth is filtered out from the results. For $D=20$km we obtain the maximum fraction of ground truth rank improvements and filtered results, respectively 12% and 4%. As $D$ increases, the fraction of affected images decreases to less than 5%. Note that the filter for $D=20$km is too aggressive, wrongly removing species from the rankings and hindering any gains from improving the ground truth’s ranking position. This effect explains why the best Top-1, Top-5 and MRR results for the filtered model are obtained for $D=40$km.
@@ 4.6 Comparative Pl@ntNet API Results
To compare Floralens with the current state-of-the-art in automatic identification of plants, and given the restrictions in accessing the APIs for most of the platforms of interest (c.f. discussion in Section  2), we opted to use Pl@ntNet for which we were generously granted free extended access. Recent work has identified Pl@ntNet as one of the best-performing platforms for this purpose [62, 43] and, therefore, a good reference point for our evaluation.
@@
The Pl@ntNet API [63] is a web service providing access to the same visual identification models used by state-of-the-art Pl@ntNet apps [6]. The API returns a set of ranked species for a given image for two models for worldwide flora: a so-called “legacy” model from 2022 (henceforth ${\rm PN}^{22}$) generated using CNN, and; a more recent model announced in July 2023 [26], generated using Vision Transformers, henceforth ${\rm PN}^{23}$. Through the API, it is also possible to filter results from the ${\rm PN}^{23}$ model so that only species occurring in a specific biogeographic region are included. One such region is Southwestern Europe including Portugal allowing the most head-to-head comparison between Floralens and Pl@ntNet possible. These results are identified by ${\rm PN}^{23\text{F}}$.
Fig 12: Pl@ntNet API: comparative MRR values.
In Figure 12, we compare the MRR values obtained for Floralens, ${\rm PN}^{22}$, ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$ for all test datasets (FLTS, PlantCLEF, and Wikipedia). We also depict separate results gauging the impact of Pl@ntNet images: the MRR for the FLTS subset excluding Pl@ntNet images (c.f. “FLTS excl. PN” series) and for the FLTS subset containing only Pl@ntNet images (c.f. “FLTS only PN” series). Recall that images from Pl@ntNet are used to train our model, which is also (of course) the case for Pl@ntNet models. In particular, part of the Pl@ntNet images we use for testing may have been used to train the Pl@ntNet models, i.e. Pl@ntNet images in our test split could be part of Pl@ntNet train splits. Hence, the set of FLTS test images excluding Pl@ntNet may in this sense be a fairer comparison between Floralens and Pl@ntNet.
Regarding the overall test datasets, we can draw the following main conclusions: (1) Floralens performs better than ${\rm PN}^{22}$ by $+0.07$ (0.75 vs. 0.68) and has almost the same performance as ${\rm PN}^{22}$ for PlantCLEF (0.73 vs. 0.72), and Wikipedia (0.72 vs 0.73); (2) Floralens performs worse than both ${\rm PN}^{23}$ and ${\rm PN}^{23\text{F}}$ by $-0.05$ for FLTS (0.75 vs. 0.80), $-0.06$ for PlantCLEF (0.73 vs. 0.79) and 0.06/0.07-0.06/-0.07- 0.06 / - 0.07 for Wikipedia (0.72 vs. 0.78 and 0.72 vs. 0.79), and (3) the Southwestern Europe species filter associated with ${\rm PN}^{23\text{F}}$ has little impact on the results, compared to ${\rm PN}^{23}$.
Focusing now on the impact of Pl@ntNet images on the FLTS results, we can see it is significant for the Pl@ntNet models, unlike Floralens which has homogeneous performance across different data sources in the FLTS as discussed previously (c.f. Figure 7(a)). It is clear that the Pl@ntNet models perform much better for Pl@ntNet images alone (“FLTS only PN” series), which amount to 35% of the entire FLTS (c.f. Table 2 in Section 3), than for the other images in the FLTS (“FLTS excl. PN” series): by +0.14 in the case of ${\rm PN}^{22}$ (0.77 vs. 0.63) and +0.06 in the case of ${\rm PN}^{23}$/${\rm PN}^{23\text{F}}$ (0.84 vs. 0.78). In contrast, the corresponding Floralens values are very close for these two sets: 0.75 and 0.74. Given Pl@ntNet models’ results seem to be biased for Pl@ntNet images (conforming to our initial hypothesis), FLTS images excluding Pl@ntNet may provide a more balanced comparative analysis. FLTS performs better than ${\rm PN}^{22}$ as discussed above (0.75 vs. 0.68), but even more so for non-Pl@ntNet images (0.75 vs. 0.63, a difference of $+0.12$); as for ${\rm PN}^{23}$/${\rm PN}^{23\text{F}}$ models, the difference is much smaller for non-Pl@ntNet images (0.75 vs 0.78, a difference of $+0.03$) than for Pl@ntNet images (0.74 vs. 0.84, a difference of $+0.10$).
(a) Results according to species growth form.
(b) Results for special categories.
Fig 13: Pl@ntNet API: comparative MRR values grouped by species growth form and special categories.
Figure 13 presents comparative MRR results again for FTLS, but grouped by (13(a)) species growth type and (13(b)) the special categories of species as analysed previously for Floralens. For a more balanced comparison, in line with the above discussion, we leave out FLTS images obtained from Pl@ntNet. The results are generally on par with the general trends already discussed. Floralens outperforms ${\rm PN}^{22}$ by a factor $+0.09$ (endemic species 13(b)) or higher up to $+0.16$ (rare species). In turn, it is outperformed by ${\rm PN}^{23}$/${\rm PN}^{23\text{F}}$ by a factor between $-0.01$ (rare species in 13(b)) and $-0.10$ (tree species in 13(a)). The cases where Floralens fares more favorably compared with ${\rm PN}^{23}$/${\rm PN}^{23\text{F}}$ are herb species regarding growth type (0.75 vs. 0.78), and rare species (0.73 vs 0.74) regarding special species’ statuses.
5 Data and Software Artifacts
@@ 5.1 Biolens Website
The Floralens model has been integrated into the Biolens project website [9]. As illustrated in the screenshots of Figure 14, the functionality is quite simple: (a) users submit photos of interest, and (b) obtain suggestions for identifications, with a textual indication of the model’s confidence – high ($>70\%$), medium ($70\%-40\%$) and low ($<40\%$). Up to 5 species are listed by the app, as long as the model outputs a minimum score of 15% for each identification. The Biolens website is hosted by a small virtual machine that requires just 2 CPU cores and 8 GB of RAM. With this configuration, invoking the model takes an average of 900 milliseconds per image, calculated directly from approximately 29,000 requests to the server between April 2021 and April 2025. As for the model itself, it takes only 8 MB of disk storage. The server configuration is lightweight as we use the TF Lite variant of the Floralens model running on the server side (similarly to the other models for other taxa also hosted on the site). The user’s browser merely displays the server’s results; it does not host the model.
@@
Fig 14: Biolens – web application screenshots.
@@ 5.2 Biolens App
Fig 15: Biolens – mobile application screenshots.
@@ 5.3 Dataset and Results
The dataset and the Python notebooks with all the code used for the results of Section 4 are available publicly from Zenodo [16]. The dataset contains the mapping between the image labels (ground truth), the image URLs from which they were retrieved, URLs for a site we maintain where all images are also stored, and GBIF identifiers when applicable (all images except those obtained from FloraOn). Ground truth and URLs are also available for the PlantCLEF and Wikipedia datasets. The Top-5 results and the corresponding confidence levels for the Floralens model and the three Pl@ntNet model variations are included for all datasets.
6 Discussion
@@
In this paper, we describe the construction of a dataset for the Portuguese flora and the derivation of a deep-learning model for the automatic identification of the species therein. The universe of species was taken from the FloraOn dataset, provided by the Sociedade Portuguesa de Botânica and compiled exclusively by specialists. The dataset was constructed based on high-quality data from several research-grade datasets available via GBIF. Besides FloraOn these include: iNaturalist, Pl@ntNet, and Observation.org. The dataset is available publicly on Zenodo [16]. The Floralens model was derived from this dataset using GAMLV which provides users with tools to derive models from datasets using off-the-shelf convolutional deep neural networks. The model is available online at the BioLens Project website and can also be used as part of the Biolens mobile application (e.g., offline, in the field).
The baseline Floralens model has good predictive power, with an AUC metric value of 0.72 and, for a reference confidence level of 0.5, values for precision and recall of 0.85 and 0.53, respectively. It also features a relatively homogeneous predictive power across all the dataset’s data sources. We measured values of 0.72 for Top-1, 0.88 for Top-5 and 0.78 for MRR. An interesting result concerns the model’s accuracy in identifying endangered, rare, protected and endemic species. Some research (e.g., [43]) suggests that current platforms struggle for these classes. The results for Floralens over these categories show that the model maintained its nominal predictive power. Using multiple images of the same specimen significantly improved the model’s accuracy for all metrics: $0.80$ for Top-1, $0.94$ for Top-5, and $0.86$ for MRR. This agrees with the findings reported by other projects (e.g., [41, 6]). Introducing filters based on species’ geographical distribution also improved, albeit less significantly, the model’s accuracy. The results testify to the accuracy and completeness of the FloraOn dataset’s geographical data. Finally, we note that the accuracy of the Floralens model is significantly better than Pl@ntnet’s legacy model (${\rm PN}^{22}$) which uses the same technology for model derivation, namely Convolutional Neural Networks. Moreover, the difference in performance between Floralens and the new Pl@ntnet models (${\rm PN}^{23}$and ${\rm PN}^{23\text{F}}$), that use Vision Transformers for model derivation, is much smaller. In this case, Pl@ntnet’s better accuracy might be attributed to using Vision Transformers. In other words, for the same model derivation technology, Floralens clearly outperforms Pl@ntnet for the Portuguese flora.
As for future work, we aim to improve the model’s accuracy by increasing the number of observations in the dataset that feature multiple images of the same specimen, eventually tagged by plant part. This will involve a concerted effort with users of the Floralens app and specialists, combined with extracting more multi-image observations from thus far unexplored publicly available datasets, e.g., Flora Incognita. Including images from Morocco and Algeria will help enhance the dataset with images from species in common with the Mediterranean flora. We also want to address limitations that arise from the inconsistent use of taxonomic names and synonyms. For example, the species listed in FloraOn as Atractylis gummifera is now known as Chamaeleon gummifer. Although not ubiquitous, it is widespread in the Mediterranean region [64]. It is not included in the Floralens dataset as there weren’t enough ($\geq$ 50) images available. However, a recent query in GBIF for Chamaeleon gummifer yields more than enough images to include the species in the Floralens dataset in a future update. Another thread worth pursuing is using Vision Transformers to derive our models. As noted in Section 2 this is a relatively recent technology when applied to computer vision problems which appears to have significant advantages over traditional approaches such as CNNs. We also plan to continue preliminary work on using image similarity models [14]. These may provide an alternative way to clinch an identification when traditional computer vision models yield low-confidence results. Hybrid classification models that combine both approaches are a possibility worth checking. Finally, more work is also required on the Biolens mobile app to improve its usability, optimize resource usage and integrate it with existing Citizen Science platforms.